OpenStudy (anonymous):

Identify the GCF 16x^4y^3 + 8x^3y^2 - 20x^2y

OpenStudy (ivettef365):

between 16 / 8 / 20 = 4 between x^4/x^3/X^2 = x^2 (use the smallest) between y^3/y^2/y = y (smallest)

OpenStudy (ivettef365):

Therefore: 4x^2y

OpenStudy (anonymous):

ok i get it now im not great with GCF thanks